OpenStudy (anonymous):

Find the sum of the infinite series if it exists. If it exists, round to one decimal place. If it does not exist, write none. 1/3+4/9+16/27+64/81

OpenStudy (tkhunny):

\(\dfrac{1}{3} + \dfrac{1}{3}\cdot\dfrac{4}{3} + \dfrac{1}{3}\cdot\left(\dfrac{4}{3}\right)^{2} + \dfrac{1}{3}\cdot\left(\dfrac{4}{3}\right)^{3} + ...\) Hmmm... I'd worry about convergence before I tried to fidn the sum.

OpenStudy (anonymous):

wait so none exists or what??

OpenStudy (tkhunny):

What do you think? A geometric series with r = 4/3 > 1.

OpenStudy (anonymous):

i dont think so but im not sure

OpenStudy (tkhunny):

|r|, the common ratio, must be less than 1 for convergence.

OpenStudy (anonymous):

ok?? im so confused

OpenStudy (tkhunny):

Confused about what? It doesn't converge. State that this is so and move on to the next problem. This one is done.